OpenStudy (anonymous):

tan^2x + cot^2x = sec^2x + cosec ^2x - 2

OpenStudy (anonymous):

tan^2x = sec^2x -1 cot^2x = csc^2x -1 ^^ Use these, you should be able to get the right side from the left.

OpenStudy (anonymous):

ohhhhh yessss :) thank you very much